I want to double-check that my method is right so please post a detailed workup

Solve the IVP = 5 3 -3 :] [:) +(64*] []-[-] Y 1
Answer #1

Please UPVOTE.

Solution Ginen I up is 63-6: 6315 (A -) za 5x+3y +6telt y= -3x+ y · 210)=1,410)= | we use elimination method to solve it. dfirst solue the homogeneous equation- 20+4x+ 440 >> D + 4D +4) * =0 =) (0+2)? 20 The roots of the auseiliarly eq are- a?+46 Giuen that the general solution is given as- alt) = Holt) + x,1t) 74(t) = (0,+Ct)elt +372 24-34 ezt 210)=1, then put to inusing initial condition y(0)=-1 3 410) = 3 + C = -3 = 3+cz = (=-6 Put C₂ =-6 in ① & @ final particular solution is- xelt) = (